Can you please help me with this system of linear equations problem
this needs to be solved by Elimination

A chemist wants to make 75 liters of 16% acid solution. He is going to make it by mixing a 10% acid and a 25% acid solution. How many liters of each will he need?
if you could just write out your work in the best way to solve this (using elimination) i'd love that
thank you once again

Answer by Alan3354(69443)

Amanda wants to make 8 gal. of a 20% saline solution by mixing together a 56% saline solution and a 8% saline solution. How much of each solution must she use?
================
e = amount of 8%
f = amount of 56%
---
e+ f = 8 (total solution)
8e + 56f = 20*8 (total saline)
---
e+ f = 8
e + 7f = 20
------------ Subtract
-6f = -12
f = 2
e = 6